OpenStudy (anonymous):

Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@312856MLP

OpenStudy (anonymous):

@XxNina26xX

OpenStudy (anonymous):

@jdorta1

OpenStudy (anonymous):

sorry i dont know.. im horrible at math =(

OpenStudy (anonymous):

its ok do you know anyone who is good at math?

OpenStudy (anonymous):

@greenlegodude57 helped me, but idk if he is online anymore or if he can help

OpenStudy (mathmale):

I need clarification of your 2.5x – 10.5 = 64(0.5x). What do you mean by 64(0.5x) ? Is this exponentiation, multiplication or what?

OpenStudy (anonymous):

It is actually 2.5x - 10.5 = 64(0.5^x)

OpenStudy (mathmale):

That info makes a huge difference. Thank you for checking and retyping your expression. You're being asked to find a solution to 2.5x - 10.5 = 64(0.5^x) by filling out a table. Let f(x)=2.5x-10.5 and let g(x)=64(0.5)^x. Choose a couple of x values and evaluate f(x) and g(x) for each. Does this result in finding an x value for which f(x)=g(x)? If so, that x value is your solution. If not, try other x values.

OpenStudy (anonymous):

would this work? f(x) = 2.5x - 10.5 g(x) = 64(0.5x) f(x) = 2.5(2) - 10.5 g(x) = 64(0.5*2) f(x) = 5 - 10.5 = -5.5 g(x) = 64(1) = 64

OpenStudy (mathmale):

Yes, but this method of finding a solution is awfully slow. Note that f(2) = -5.5 is not equal to g(2) = 64. In your shoes I'd try an alternative approach: Graph f(x)=2.5x-10.5 and g(x) on the same set of axes and determine where the two graphs intersect. That will give you the "solution" to this system of equations. I was able to this quickly on my TI-84 calculator. Try that, if you have one. Also, please note that if you choose x=1, g(1)=64(0.5)^1 = 64(0.5) = 32.